OpenStudy (darkbluechocobo):

Which graph is f(x) = -(x + 1)^3 − 2

OpenStudy (darkbluechocobo):

so I limited it down to two

OpenStudy (darkbluechocobo):

but I dont know if it flips the way its going or not

jimthompson5910 (jim_thompson5910):

when x = 0, f(x) = ???

OpenStudy (darkbluechocobo):

0?

jimthompson5910 (jim_thompson5910):

f(x) = -(x + 1)^3 − 2 f(0) = -(0 + 1)^3 − 2 ... replace EVERY x with 0 f(0) = ???

OpenStudy (darkbluechocobo):

f(0)=-1

jimthompson5910 (jim_thompson5910):

f(x) = -(x + 1)^3 − 2 f(0) = -(0 + 1)^3 − 2 ... replace EVERY x with 0 f(0) = -(1)^3 - 2 f(0) = -1 - 2 f(0) = -3

jimthompson5910 (jim_thompson5910):

because f(0) = -3 we know that (0,-3) is a point on the curve

OpenStudy (anonymous):

Here is a pic for you.